Improper Integrals (Integrals with Infinite Limits)


∫dt/t^3 from1 to ∞


1
Expert's answer
2021-05-07T10:21:18-0400

"\int_1^\infin\dfrac{1}{t³}dt = \int_1^\infin t^{-3}dt"


"= [\dfrac{t^{-2}}{-2}]^\infin_1 = [-0.5t^{-2}]^\infin_1"


"= 0-(-0.5(1^{-2}) = 0.5"